OpenStudy (anonymous):

Rate of Change Problem!! Please help :/ At noon, shop A is 100km west of ship B. Ship A is sailing south at 35 km/h and ship B is sailing north at 25 km/h. How fast is the distance between ships changing at 4:00 PM?

OpenStudy (anonymous):

i assumed that the noon mean 1200pm, and i assume that u want the changes of distance between both ship: distance = sqrt((horizontal length)²+(vertical length)²) = (100²+(60( (35+25) * t )²)^1/2 (35+25) is from the velocity of both ship, becuz they re moving apart from each other. t is the time velocity or speed=(displacement or distance)/time , if u use 'time' times velocity, u will get the distance. d/dx (100²+(60( (35+25) * t )²)^1/2 =(3600t)/(100²+(60t²))^1/2 i hope thats right